# MLVU: A Comprehensive Benchmark for Multi-Task Long Video Understanding

### v1

@@ Abstract
The evaluation of Long Video Understanding (LVU) performance poses an important but challenging research problem. Despite previous efforts, the existing video understanding benchmarks are severely constrained by several issues, especially the insufficient lengths of videos, a lack of diversity in video types and evaluation tasks, and the inappropriateness for evaluating LVU performances. To address the above problems, we propose a new benchmark, called MLVU [^1] MLVU Repository: https://github.com/FlagOpen/FlagEmbedding. (Multi-task Long Video Understanding Benchmark), for the comprehensive and in-depth evaluation of LVU. MLVU presents the following critical values: 1) The substantial and flexible extension of video lengths, which enables the benchmark to evaluate LVU performance across a wide range of durations. 2) The inclusion of various video genres, e.g., movies, surveillance footage, egocentric videos, cartoons, game videos, etc., which reflects the models’ LVU performances in different scenarios. 3) The development of diversified evaluation tasks, which enables a comprehensive examination of MLLMs’ key abilities in long-video understanding. The empirical study with 20 latest MLLMs reveals significant room for improvement in today’s technique, as all existing methods struggle with most of the evaluation tasks and exhibit severe performance degradation when handling longer videos. Additionally, it suggests that factors such as context length, image-understanding quality, and the choice of LLM backbone can play critical roles in future advancements. We anticipate that MLVU will advance the research of long video understanding by providing a comprehensive and in-depth analysis of MLLMs.
@@ 1 Introduction
Large language models (LLMs) are growing into a general solution for numerous AI tasks [6; 39]. In recent years, it becomes increasingly emphasized to extend LLMs with multi-modal capabilities and thus bring the Multi-modal LLM, namely, MLLM. Remarkably, it has been made possible for today’s MLLMs to perceive information in texts, images, videos, etc., and solve complicated problems in physical environments [1; 38]. Along with the development of MLLMs, new benchmarks are continuously created to facilitate comprehensive and in-depth analysis of MLLMs [50; 27; 9; 21].
However, it remains a great challenge to evaluate the MLLMs’ long-video understanding (LVU) performances given the following limitations. Firstly, the majority of existing video understanding benchmarks are made up of short videos [47; 21; 19; 31; 15], whose lengths can be merely a few seconds. As a result, they are insufficient to reflect the MLLMs’ long-video understanding capabilities. Secondly, there is a severe lack of diversity for the existing LVU benchmarks in terms of video genres and evaluation tasks. On one hand, the existing benchmarks are usually created based on one type of videos, e.g., movies [35; 22]. On the other hand, it may only contain one specific type of task in each individual benchmark, for example, the captioning task in [47], the temporal perception task in [49], and the action understanding task in [41; 11]. Due to the above limitations, existing benchmarks have difficulty in performing a comprehensive evaluation of LVU performances. Last but not least, many previous evaluation tasks are not properly designed for LVU, as they can be solved without using the complex information from long videos. For example, many questions are simply about one single frame in the long videos [35]. Besides, numerous others are about popular movies and celebrities [22], which can be answered directly by MLLMs based on the textual prompts.
@@ 2 Related Work
Multimodal large language models (MLLMs) have gained wide interests from both academic and industrial sectors. Recent studies have made significant progresses in this field by integrating LLM backbones with visual encoders and adapters, and fine-tuning the entire architecture through visual instruction [25; 55; 7]. Based on the same philosophy, MLLMs have been further developed for video processing using video instruction datasets and specialized video adapters [52; 28; 20; 48; 23; 21]. However, most existing models are optimized for short videos, typically under one minute, due to the difficulty in establishing sufficient context for longer videos. To address this challenge, researchers have explored compact representations of videos. For example, LLaMa-Vid [22] compresses each video frame into two tokens, allowing the model to handle videos several hours long. Methods like MovieChat [35] and MA-LMM [12] introduce specialized memory components for recursive video processing. Additionally, it is also explored to make selective usage of frames or clips from long videos based on retrievers or agents [46; 33; 40]. Despite these progresses, it remains an open problem for MLLMs to effectively handle long videos.
@@
With the unprecedented interest in MLLMs, the creation of benchmarks for these models has become increasingly emphasized (as advanced by MMMU [50], MME [9], and many other pioneering works). In video understanding, the research community has made significant efforts as well, particularly for short videos. There are specialized benchmarks for temporal perception [49; 42], action understanding [42; 41], video classification [14], video reasoning [45; 44], and video captioning [47; 30]. Recently, MVBench [21] provides a comprehensive short-video benchmark to evaluate general capabilities via question-answering. For long video understanding, people seek to leverage long-form videos, like movies, to create benchmarks. For example, the LVU dataset [43] presents movie understanding tasks, such as predicting release years and identifying character relationships. Similarly, LLaMA-Vid [22] created a movie question-answering dataset based on MovieNet [13]. Despite using long videos, many questions are about common-sense knowledge of the movies. As a result, they could be directly answered without using the video’s information. In contrast, MovieChat [35] makes use of diversified videos and avoids specific character names or plot details in its questions. Considering that many of the questions target on exact time segments, it potentially degrades the tasks to short-video or image understanding problems. Beyond movies, there are other task-specific benchmarks like EgoSchema [29], which presents video reasoning tasks using first-person footage from Ego4D [11]. However, these specialized benchmarks only focus on one single aspect of MLLMs, rather than offering a comprehensive analysis of long video understanding. Therefore, it remains essential to develop a comprehensive benchmark with carefully designed tasks to effectively evaluate MLLMs’ capabilities in understanding long videos.
@@ 3.1 Overview
MLVU offers a comprehensive collection of videos of different categories (Figure 1 Middle). There are typical real-world videos, which includes movies [51; 35], documentaries [35], TV series, egocentric videos [11], and surveillance footage [37]. Meanwhile, there are also important simulated videos from animated series and game videos [8].
@@ 4.2 Overall Performance
On the other hand, although GPT-4o achieves a huge advantage in comparison with other methods, it actually struggles to handle most of the tasks in the benchmark. For example, it only ends up with 64.8% in dealing with the needle question-answering (NQA) task. Whereas its analogue tasks in the text domain, e.g., NIHS (Needle-In-the-HayStack-Search) and Passkey Retrieval, can be effectively conquered by many of the existing long LLMs [10; 53]. At the same time, it exhibits even less reliability given tasks like ego-reasoning (ER), action ordering (AO), and action count (AC), while other baseline methods produce even worse performances in these scenarios. The above observations indicate that long-video understanding remains a tough challenge for today’s MLLMs.
@@ H.2 Comprehensive evaluation results on MLVU Time-ladder
MLVU can be viewed and downloaded on GitHub at https://github.com/FlagOpen/FlagEmbedding/tree/master/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU. We assure its long-term preservation for future reference and use. The annotations for questions and answers are provided in the JSON file format, while the raw videos are available in the MP4 format.
@@
Answer: The benchmark is available on GitHub at https://github.com/FlagOpen/FlagEmbedding/tree/master/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.

### v2

@@ Abstract
The evaluation of Long Video Understanding (LVU) performance poses an important but challenging research problem. Despite previous efforts, the existing video understanding benchmarks are severely constrained by several issues, especially the insufficient lengths of videos, a lack of diversity in video types and evaluation tasks, and the inappropriateness for evaluating LVU performances. To address the above problems, we propose a new benchmark, called MLVU [^1] MLVU Repository: https://github.com/JUNJIE99/MLVU. (Multi-task Long Video Understanding Benchmark), for the comprehensive and in-depth evaluation of LVU. MLVU presents the following critical values: 1) The substantial and flexible extension of video lengths, which enables the benchmark to evaluate LVU performance across a wide range of durations. 2) The inclusion of various video genres, e.g., movies, surveillance footage, egocentric videos, cartoons, game videos, etc., which reflects the models’ LVU performances in different scenarios. 3) The development of diversified evaluation tasks, which enables a comprehensive examination of MLLMs’ key abilities in long-video understanding. The empirical study with 20 latest MLLMs reveals significant room for improvement in today’s technique, as all existing methods struggle with most of the evaluation tasks and exhibit severe performance degradation when handling longer videos. Additionally, it suggests that factors such as context length, image-understanding quality, and the choice of LLM backbone can play critical roles in future advancements. We anticipate that MLVU will advance the research of long video understanding by providing a comprehensive and in-depth analysis of MLLMs.
@@ 1 Introduction
Large language models (LLMs) are growing into a general solution for numerous AI tasks [6; 40]. In recent years, it becomes increasingly emphasized to extend LLMs with multi-modal capabilities and thus bring the Multi-modal LLM, namely, MLLM. Remarkably, it has been made possible for today’s MLLMs to perceive information in texts, images, videos, etc., and solve complicated problems in physical environments [1; 39]. Along with the development of MLLMs, new benchmarks are continuously created to facilitate comprehensive and in-depth analysis of MLLMs [51; 27; 9; 21].
However, it remains a great challenge to evaluate the MLLMs’ long-video understanding (LVU) performances given the following limitations. Firstly, the majority of existing video understanding benchmarks are made up of short videos [48; 21; 19; 32; 15], whose lengths can be merely a few seconds. As a result, they are insufficient to reflect the MLLMs’ long-video understanding capabilities. Secondly, there is a severe lack of diversity for the existing LVU benchmarks in terms of video genres and evaluation tasks. On one hand, the existing benchmarks are usually created based on one type of videos, e.g., movies [36; 22]. On the other hand, it may only contain one specific type of task in each individual benchmark, for example, the captioning task in [48], the temporal perception task in [50], and the action understanding task in [42; 11]. Due to the above limitations, existing benchmarks have difficulty in performing a comprehensive evaluation of LVU performances. Last but not least, many previous evaluation tasks are not properly designed for LVU, as they can be solved without using the complex information from long videos. For example, many questions are simply about one single frame in the long videos [36]. Besides, numerous others are about popular movies and celebrities [22], which can be answered directly by MLLMs based on the textual prompts.
@@ 2 Related Work
Multimodal large language models (MLLMs) have gained wide interests from both academic and industrial sectors. Recent studies have made significant progresses in this field by integrating LLM backbones with visual encoders and adapters, and fine-tuning the entire architecture through visual instruction [25; 56; 7]. Based on the same philosophy, MLLMs have been further developed for video processing using video instruction datasets and specialized video adapters [53; 29; 20; 49; 23; 21]. However, most existing models are optimized for short videos, typically under one minute, due to the difficulty in establishing sufficient context for longer videos. To address this challenge, researchers have explored compact representations of videos. For example, LLaMa-Vid [22] compresses each video frame into two tokens, allowing the model to handle videos several hours long. Methods like MovieChat [36] and MA-LMM [12] introduce specialized memory components for recursive video processing. Additionally, it is also explored to make selective usage of frames or clips from long videos based on retrievers or agents [47; 34; 41]. Despite these progresses, it remains an open problem for MLLMs to effectively handle long videos.
@@
With the unprecedented interest in MLLMs, the creation of benchmarks for these models has become increasingly emphasized (as advanced by MMMU [51], MME [9], and many other pioneering works). In video understanding, the research community has made significant efforts as well, particularly for short videos. There are specialized benchmarks for temporal perception [50; 43; 28], action understanding [43; 42], video classification [14], video reasoning [46; 45], and video captioning [48; 31]. Recently, MVBench [21] provides a comprehensive short-video benchmark to evaluate general capabilities via question-answering. For long video understanding, people seek to leverage long-form videos, like movies, to create benchmarks. For example, the LVU dataset [44] presents movie understanding tasks, such as predicting release years and identifying character relationships. Similarly, LLaMA-Vid [22] created a movie question-answering dataset based on MovieNet [13]. Despite using long videos, many questions are about common-sense knowledge of the movies. As a result, they could be directly answered without using the video’s information. In contrast, MovieChat [36] makes use of diversified videos and avoids specific character names or plot details in its questions. Considering that many of the questions target on exact time segments, it potentially degrades the tasks to short-video or image understanding problems. Beyond movies, there are other task-specific benchmarks like EgoSchema [30], which presents video reasoning tasks using first-person footage from Ego4D [11]. However, these specialized benchmarks only focus on one single aspect of MLLMs, rather than offering a comprehensive analysis of long video understanding. Therefore, it remains essential to develop a comprehensive benchmark with carefully designed tasks to effectively evaluate MLLMs’ capabilities in understanding long videos.
@@ 3.1 Overview
MLVU offers a comprehensive collection of videos of different categories (Figure 1 Middle). There are typical real-world videos, which includes movies [52; 36], documentaries [36], TV series, egocentric videos [11], and surveillance footage [38]. Meanwhile, there are also important simulated videos from animated series and game videos [8].
@@ 4.2 Overall Performance
On the other hand, although GPT-4o achieves a huge advantage in comparison with other methods, it actually struggles to handle most of the tasks in the benchmark. For example, it only ends up with 64.8% in dealing with the needle question-answering (NQA) task. Whereas its analogue tasks in the text domain, e.g., NIHS (Needle-In-the-HayStack-Search) and Passkey Retrieval, can be effectively conquered by many of the existing long LLMs [10; 54]. At the same time, it exhibits even less reliability given tasks like ego-reasoning (ER), action ordering (AO), and action count (AC), while other baseline methods produce even worse performances in these scenarios. The above observations indicate that long-video understanding remains a tough challenge for today’s MLLMs.
@@ H.2 Comprehensive evaluation results on MLVU Time-ladder
MLVU can be viewed and downloaded on GitHub at https://github.com/JUNJIE99/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU. We assure its long-term preservation for future reference and use. The annotations for questions and answers are provided in the JSON file format, while the raw videos are available in the MP4 format.
@@
Answer: The benchmark is available on GitHub at https://github.com/JUNJIE99/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.
@@
Answer: Yes. We will post new update in https://github.com/JUNJIE99/MLVU and https://huggingface.co/datasets/MLVU/MVLU if there is any.
@@
Answer: Yes. Old versions will also be hosted in https://github.com/JUNJIE99/MLVU and https://huggingface.co/datasets/MLVU/MVLU.